OpenStudy (anonymous):

i have x+2=(1/2)y=z-3 and how do i get v=<1,2,1> ?

OpenStudy (anonymous):

can somebody help me?

OpenStudy (anonymous):

x+2=(1/2)y=z-3 and its vector is v=<1,2,1>

OpenStudy (anonymous):

how do i get Vector?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

v=<at,bt,ct>

OpenStudy (zzr0ck3r):

im confused by the question...what does v have to do with the first equation?

OpenStudy (anonymous):

how do i get v=<1,2,1> from that equation x+2=(1/2)y=z-3

OpenStudy (anonymous):

okay, x+2=(1/2)y=z-3 what's the v=<at,bt,ct> ?

OpenStudy (anonymous):

i know x y z should be equal obviously. if you put 1,2,1 they all equal to 1

OpenStudy (anonymous):

can't i make them equal 0 or 100 or something? its vector should be still the same?